# Delay-Multiply-And-Sum Beamforming for Real-Time In-Air Acoustic Imaging

### v1

@@ Abstract
In-air acoustic imaging systems demand beamforming techniques that offer a high dynamic range and spatial resolution while also remaining robust. Conventional Delay-and-Sum (DAS) beamforming fails to meet these quality demands due to high sidelobes, a wide main lobe and the resulting low contrast, whereas advanced adaptive methods are typically precluded by the computational cost and the single-snapshot constraint of real-time field operation. To overcome this trade-off, we propose and detail the implementation of higher-order non-linear beamforming methods using the Delay-Multiply-and-Sum technique, coupled with Coherence Factor weighting, specifically adapted for ultrasonic in-air microphone arrays. Our efficient implementation allows for enabling GPU-accelerated, real-time performance on embedded computing platforms. Through validation against the DAS baseline using simulated and real-world acoustic data, we demonstrate that the proposed method provides significant improvements in image contrast, establishing higher-order non-linear beamforming as a practical, high-performance solution for in-air acoustic imaging.
@@ I Introduction
In-air acoustic imaging is a technique in which an array of microphones is used to reconstruct the incoming wave field that is being emitted by one or more sources in the environment. The fundamental goal of this process is to convert the time or frequency-domain signals recorded across an array of microphones into a spatially localized map that visualizes the power distribution. This technique has a wide range of applications, one of which is the detection of pressurized gas or air leaks in industrial environments [guentherAutomatedDetectionCompressed2016, steckelUltrasoundbasedAirLeak2014]. High-pressure gas escaping through a small orifice generates acoustic energy due to the turbulent, high-velocity flow. The ability to precisely localize these leaks in complex piping systems is vital for energy efficiency, environmental compliance, and worker safety. Furthermore, non-contact fault detection and condition monitoring of rotating machinery is another major driver [alousifMachineryFaultDetection2021, dadoucheSensitivityAirCoupledUltrasound2008, steckelToolWearPrediction2024, verellenBeamformingAppliedUltrasound2021]. Mechanical faults, such as bearing wear produce ultrasonic emissions long before they generate significant audible noise or heat, making acoustic imaging a viable tool for predictive maintenance. It is also used in aeroacoustics for the analysis of wing performance [merino-martinezReviewAcousticImaging2019]. Finally, ultrasonic imaging is of importance for in-air sonar sensing[allevatoEmbeddedAircoupledUltrasonic2020, jansenStabilizedAdaptiveSteering2024, moto:c:irua:166448_kers_comp]. It is used as a pulse-echo sensor in for example (3D) robotic navigation and mapping. It allows for robust depth sensing and is can be of critical value for autonomous systems operating in visually obscured environments (by dust, rain, fog, …) where other sensors such as camera’s may fail.
@@
Figure 1: On the left side of the figure we depicted a 3D in-air sonar device (eRTIS[laurijssen2025ruggedizedultrasoundsensingharsh]) as used in the experiments with the typical situation of an active pulse-echo measurement. The signal from the emitter reflects on an object and is received by all microphones, which then goes through the several signal processing steps to result in the final acoustic image. The steps used in this processing pipeline are matched filtering (pulse compression), beamforming and a final envelope extraction phase. In this paper, we will focus on improvements of the beamforming stage, which converts the signals of the microphone array into a spatialized set of signals. The figure furthermore shows the coordinate system with the azimuth angle $\theta$ and elevation angle $\varphi$, used to define directions of interests $\psi$ or sound locations. In the right sketch of the acoustic image, the vertical axis uses the range dimension $r$, converted from the time domain $t$.
@@
A general recurrence in these applications that make use of acoustic imaging techniques is the usage of of spatial signal processing. This implies the usage of some kind of sensor array, and an algorithm that converts the multitude of input signals into some kind of spatially localized signal. Array signal processing has known a very long and rich history, and several seminal books and papers have been written on this subject [treesOptimumArrayProcessing2002, krimTwoDecadesArray1996, pesaventoThreeMoreDecades2023]. When taking a zoomed-out view, the domain can be simplified and categorized into the following categories:
  • Spatial filtering: Delay-and-Sum (DAS) [krimTwoDecadesArray1996, van1988beamforming]
@@
  • Super-resolution filtering : Minimum Variance Distortionless Response (MVDR) [treesOptimumArrayProcessing2002, moto:c:irua:172542_vere_ultr]
  • Spatial spectrum estimator: Multiple Signal Classification (MUSIC) [moto:c:irua:166448_kers_comp, labyed2013super]
  • Compressed sensing: Sparse representations [steckelSparseDecompositionInair2014, sonIrregularMicrophoneArray2022]
  • Machine-learning based methods [steckel2024echopt2, lafontaine2023denoising, mansour2018sparse]
@@
The core differences between these listed techniques lie in how they approach the fundamental trade-off between image quality characteristics such as spatial resolution, sidelobe level, dynamic range, and robustness to noise as well as to correlated sources. Furthermore, computational cost has to be taken into account for practical, real-world systems. While advanced methods (like MVDR and MUSIC) offer superior resolution compared to conventional DAS, they often incur high computational complexity, and rely on assumptions that can degrade real-time performance and dynamic range in complex, real-world ultrasonic environments. More specifically, most of the so-called super-resolution techniques use spatial statistics of the incoming wavefield, requiring multiple measurements (snapshots) for the estimation thereof. In acoustic imaging, the luxury of multiple snapshots is not available, as this would incur a measurement time that is too large given the slow speed of sound[moto:c:irua:172542_vere_ultr].
Given the constraints of real-time, high dynamic range acoustic imaging, there is a need for a spatial signal processing algorithm that offers better image quality than for example DAS, without incurring the computational complexity and model dependence of MVDR or MUSIC.
@@
In this paper, we take inspiration from medical ultrasonic imaging, a field where contrast and resolution are paramount, from their usage of non-linear beamforming methods, which have been shown to drastically improve image quality. This paper focuses on the application of the Delay-Multiply-and-Sum (DMAS) combined with the optional Coherence Factor (CF) weighting for in-air acoustic imaging. The core principle of DMAS is to introduce a non-linear multiplication of delayed (or so-called pre-steered or pre-aligned) microphone signals. This operation effectively suppresses side-lobes and enhances image contrast by exploiting spatial coherence. Meanwhile, the coherence factor is used as a spatial weighting function to further reduce imaging artifacts. This methodology has demonstrated superior contrast ratios and increased range resolution in biomedical applications [lim2008confocal, dmasmatrone, mulani2022higher, Mantalenadmasius2025] but remains relatively unexplored in the context of in-air acoustic arrays.
@@
Figure 2: Directional component of the point-spread function in both azimuth and elevation, plotted using a Lambert Equal Area Projection. The resulting PSF shows an higher dynamic range for increasing orders of DMAS processing. Furthermore, the addition of the CF post-processing step further increases the dynamic range of the PSF. Gridlines are spaced by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$.
@@ II Methods
This section details the mathematical and computational methodology used to generate an acoustic image and perform source localization from raw microphone signals. The contribution of this paper focuses on the beamforming step of the imaging process, which will be reflected in this section as well. However, the overall process, as implemented in our general purpose signal processing pipeline[jansen2019gpu], follows a sequential order with additional optional pre- and post-processing steps and filters, also visualized in Fig. 1:
@@
The system consists of an array of $N$ elements, in our case microphones. The received signal at the $i$-th microphone is denoted as $m_{i}(t)$. To form an image, a set of $M$ directions of interest is defined. For each direction $\psi$, a time delay $\tau_{i,\psi}$ is calculated for each $i$-th microphone, representing the propagation time of an acoustic wave from the direction to the $i$-th microphone. In the case of 3D beamforming, these directions are defined by their azimuth and elevation angles $(\theta,\varphi)$. Fig. 1 visualizes the coordinate system and the signal processing pipeline. The delays for these directions are pre-computed and stored in a delay matrix look-up table. The pre-steering operation is to time-shift the received signals according to these delays, effectively compensating them. For a given time $t$ and direction $\psi$, we define the delayed signal $x_{i}(t,\psi)$ as:
@@ II-A Delay-and-Sum (DAS) Beamforming
DAS, as a non-adaptive beamformer, treats desired signals and noise/clutter identically, leading to high sidelobes and low image resolution and contrast, which is especially problematic in the noisy, attenuating environment of for in-air acoustic imaging.
@@ II-B Delay-Multiply-and-Sum (DMAS) Beamforming
DMAS is a non-linear technique that involves multiplying combinations of delayed signals before summation [lim2008confocal, dmasmatrone]. It is generally described by:
@@
The multiplication will cause lower sidelobes and a higher dynamic range by acting as a correlation function, which in effect suppresses low-coherence parts in the image signals. As these low-coherence parts are typically sidelobes in the image, and not real reflected signals, this operation increases image quality. High-order DMAS methods have been proposed to further suppress more coherent artifacts [mulani2022higher]. Let the $n$-th signed root of the delayed signal $x_{i}$ be:
@@
Using this $n$-th root signal $s_{i}^{(n)}$, the higher-order version of DMAS for the $n$-th order is given by:
@@ II-C Efficient Higher-Order DMAS Beamforming
To achieve an efficient implementation, we use the Newton-Girard identities [macdonald1998symmetric], which can provide an explicit formula for $E_{n}$ in terms of the power sums of its arguments. The first step is to compute the necessary power sums, $P_{k}^{(n)}$, for $k=1,\ldots,n$:
@@
Note that order 2 is equivalent to DMAS as in Equation (3). For any given order $n$, our method involves first computing the power sums $P_{1}^{(n)},\ldots,P_{n}^{(n)}$ in $\mathrm{O}(N)$ time, and then substituting these values into the corresponding explicit formula above to obtain $S_{\textsl{DMAS}}^{(n)}=E_{n}$. This non-recursive, explicit method is highly efficient for a fixed set of orders.
@@ II-D Coherence Factor (CF) Weighting
The Coherence Factor (CF) is an optional spatial weighting applied to the beamformer output to suppress incoherent noise and sidelobes [coherencehollman1999]. It measures the spatial coherence of the delayed signals across the array aperture, and is defined as the ratio of the coherent power to the incoherent power. The specific formulation we implemented can be written as follows:
@@
II-E Computational GPU Acceleration
Even with the proposed efficient implementation, the entire signal processing is computationally demanding, especially for real-time processing on embedded or edge-computing devices. To achieve high performance, we have implemented the necessary beamforming operations on GPU hardware. We implemented a kernel function that achieves parallelism by launching a separate thread for each output pixel (i.e., for each $(t,\psi)$). The efficient $\mathrm{O}(N)$ DMAS formulation is critical, as it allows each thread to compute its output by iterating only once over the $N$ microphones. Our implementation is written in NVIDIA’s CUDA framework, and the source code for this can be shared upon request.
@@ III Experimental Results And Validation
To evaluate the performance and merit of the DMAS algorithms for in-air sensing, we did several experiments. We want to note that in our results, we restrict our comparison solely to DAS against our proposed method because we want to focus on real-world conditions for a broad field of realistic application scenarios. This means we want to operate using only single snapshot, precluding the use of adaptive methods like MVDR which require multiple snapshots to accurately estimate data statistics. Furthermore, we require a method that functions effectively with any arbitrary microphone array configuration. Indeed, single-snapshot adaptations for super-resolution techniques exists, but require the application of spatial smoothing, which puts constraints on the array layout[moto:c:irua:172542_vere_ultr]. Finally, the imaging systems we want to apply these techniques to are inherently broadband, with typical frequency ranges from $25\text{\,}\mathrm{kHz}$ to $50\text{\,}\mathrm{kHz}$. For these wideband applications, techniques like DAS and DMAS inherently function well as they are time-domain processing algorithms, unlike some frequency-domain or narrowband adaptive techniques such as MVDR or MUSIC.
@@ III-A Simulation Analysis
In order to quantify the performance of DMAS-class of beamformers in in-air sonar applications, we did several simulations of key metrics of this imaging system. As a first validation, we calculated the Point-Spread Function (PSF) of the whole imaging system. A PSF is a commonly used characterization of an imaging system, and describes the response of the system to a Dirac point source. In an active imaging system the point-spread function consists of three dimensions: two direction dimensions and one range dimension. In Fig. 2 we have plotted the directional part of the PSF in the noise-free case. From this figure it becomes apparent that the dynamic range of the point-spread function becomes much greater for increasing DMAS orders, reaching almost $80\text{\,}\mathrm{dB}$ in the DMAS5 class. Furthermore, the CF post-processing step further increases the dynamic range of the PSF. Little effect can be observed on the width of the main-lobe of the PSF.
A second view of the point-spread function can be made in the range/direction plane, more specifically in the horizontal plane for an elevation of $0\text{\,}\mathrm{\SIUnitSymbolDegree}$. Effectively, the PSF is sliced in the horizontal plane, and shown in a range/azimuth plot. The resulting ensemble of these PSFs can be found in Fig. 3. The dynamic range of the PSF improves for increasing orders of DMAS processing, with an additional gain in the dynamic range when applying the CF post-processing step. As expected, the range resolution is not impacted by the DMAS algorithms, because of to the separability of the range and direction dimensions in far-field broadband imaging systems, highlighting the function of DMAS as a pure beamforming operation.
@@
The literature of medical imaging does highlight increased range resolution when applying DMAS, which can be explained by the fact that in medical imaging, the imaging takes place in the near-field regime (with the sensor array not being considered ”small” in respect to the imaging scene). Therefore, range-focusing through beamforming is possible, which causes an effect on the range resolution of the DMAS operations.
As with in-air sonar sensing we typically operate in a pure far field regime (with array apertures of $10\text{\,}\mathrm{cm}$ diameter, and imaging ranges spanning up from $5\text{\,}\mathrm{m}$ to $10\text{\,}\mathrm{m}$), the separability of the range and directional imaging steps holds, and therefore, the lack of increase in range resolution when applying DMAS is expected.
In a second experiment, we calculated the signal-to-noise ratio (SNR) of the acoustic image for a varying SNR of the input microphone signals, using the following procedure: Let the amplitude of the echo signal for a source placed at broadside of the array at $1\text{\,}\mathrm{m}$ range in the microphone channels $m_{i}(t)$ be equal to 1.
We add a white gaussian noise signal $n_{i}(t)$ to the microphone signals in the following way:
@@
We calculated these image SNR quantities for an input range of $\mathrm{mic}$ ranging from -$40\text{\,}\mathrm{dB}$ to $10\text{\,}\mathrm{dB}$. The results of these simulations can be found in Fig. 4, with the panels showing the SNR results for the imaging with and without coherence factor post-processing on the left and right respectively.
From these plots the superior noise suppression of the DMAS-class of algorithms can be observed, with increasing image SNR values for increasing DMAS orders. The addition of the CF post-processing step further increases the image SNR over all algorithms.
As a final quantification experiment, we measured the spatial resolution of the DMAS-class beamformers. For this, we implemented a simulation in which two reflectors were spaced at the same range ($1.5\text{\,}\mathrm{m}$) and are placed symmetrically with respect to the X-axis with a half-angle $a$, as can be seen in Fig. 5 on the bottom right corner. We then calculated the acoustic image $I$ in the horizontal plane, and selected the range slice corresponding to the reflector range. We then aggregated the angular responses for all inner-angle combinations of reflectors in the scene, which are the results shown in Fig. 5. From these plots it becomes clear that the main benefit of the DMAS methods is an increase in dynamic range, but not in spatial resolution, which remains mostly constant over the spacing of the targets. This confirms the hypothesis that the spatial resolution mainly originates from the layout of the microphone array, and cannot be drastically improved by applying the DMAS-class algorithms for beamforming.
][!ht] Refer to caption
@@
To further support the claim that the spatial resolution of the imaging system only marginally depends on the order of the DMAS beamformers, we calculated the beamwidth of the various beamformer implementations for varying array sizes. We simulated circular arrays with uniformly spaced microphones inside of the array, and varied the radius of the arrays from $1\text{\,}\mathrm{cm}$ to $6\text{\,}\mathrm{cm}$, for which the microphone count varied from 19 to 513 microphones (as we kept the spacing between the microphones constant on an hexagonal lattice with an edge length of $5\text{\,}\mathrm{mm}$). For each of the array shapes we then calculated the point-spread function in the horizontal plane and calculated the -$3\text{\,}\mathrm{dB}$ beamwidth of the PSF. This does not cause a loss of generality, as the PSF of a circularly symmetric array is circularly symmetric as well.
We performed this calculation for all beamformer implementations, and plotted the resulting beamwidths in function of the array radius in Fig. 6. The right panel details the experiment and illustrated the more narrow PSF as the array size increases. Both graphs show that the major decrease in beamwidth is caused by the array shape, and less so for the various beamformer implementations, with the beamwidths converging for the larger (yet still very practical) array sizes. This convergence of spatial resolution for larger arrays coincides with the findings depicted in Fig. 5, where it could be observed that the spatial resolution is not greatly increased using the DMAS-class of beamformers.

### v2

@@ Abstract
In-air acoustic imaging systems demand beamforming techniques that offer high dynamic range and spatial resolution while remaining robust, yet conventional Delay-and-Sum (DAS) beamforming suffers from high sidelobes and low contrast. Conversely, advanced adaptive methods are typically too computationally expensive for real-time field operation. To overcome this trade-off, we propose a higher-order non-linear beamforming method using the Delay-Multiply-and-Sum (DMAS) technique with Coherence Factor weighting, specifically adapted for in-air microphone arrays. By leveraging Newton-Girard binomial expansion, we reduce the complexity of $n$-th order DMAS to $\mathrm{O}(N)$, enabling GPU-accelerated, real-time performance on embedded platforms, with the implementation also made available as open-source. Through validation against simulated and real-world data, we demonstrate that the proposed 5th-order DMAS with CF achieves a dynamic range improvement of nearly $40\text{ dB}$ over DAS, alongside significantly higher image SNR and artifact suppression. We demonstrate real-time viability on various GPU platforms, establishing this approach as a practical, high-performance solution for 3D in-air sonar and industrial acoustic imaging.
@@ I Introduction
In-air acoustic imaging is a technique in which an array of microphones is used to reconstruct the incoming wave field that is being emitted by one or more sources in the environment. The fundamental goal of this process is to convert the time or frequency-domain signals recorded across an array of microphones into a spatially localized map that visualizes the power distribution.
@@
This technique has a wide range of applications, one of which is the detection of pressurized gas or air leaks in industrial environments [4, 30]. High-pressure gas escaping through a small orifice generates acoustic energy due to the turbulent, high-velocity flow. The ability to precisely localize these leaks in complex piping systems is vital for energy efficiency, environmental compliance, and worker safety. Furthermore, non-contact fault detection and condition monitoring of rotating machinery is another major driver [2, 3, 27, 34]. Mechanical faults, such as bearing wear produce ultrasonic emissions long before they generate significant audible noise or heat, making acoustic imaging a viable tool for predictive maintenance. It is also used in aeroacoustics for the analysis of wing performance [18]. Finally, ultrasonic imaging is of importance for in-air sonar sensing[1, 8, 9]. It is used as a pulse-echo sensor in, for example, (3D) robotic navigation and mapping. It allows for robust depth sensing and can be of critical value for autonomous systems operating in visually obscured environments (by dust, rain, fog, etc.) where other sensors such as cameras may fail.
@@
Figure 1: On the left side of the figure we depict a 3D in-air sonar device (eRTIS[13]) as used in the experiments with the typical situation of an active pulse-echo measurement. The signal from the emitter reflects on an object and is received by all microphones, which then goes through the several signal processing steps to result in the final acoustic image. The steps used in this processing pipeline are matched filtering (pulse compression), beamforming and a final envelope extraction phase. In this paper, we will focus on improvements of the beamforming stage, which converts the signals of the microphone array into a spatialized set of signals. The figure furthermore shows the coordinate system with the azimuth angle $\theta$ and elevation angle $\varphi$, used to define directions of interest $\psi$ or sound locations. In the right sketch of the acoustic image, the vertical axis uses the range dimension $r$, converted from the time domain $t$.
@@
A general recurrence in these applications that make use of acoustic imaging techniques is the usage of spatial signal processing. This implies the usage of some kind of sensor array, and an algorithm that converts the multitude of input signals into some kind of spatially localized signal. Array signal processing has a very long and rich history, and several seminal books and papers have been written on this subject [31, 10, 21]. When taking a zoomed-out view, the domain can be simplified and categorized into the following categories:
  • Spatial filtering: Delay-and-Sum (DAS) [10, 32]
@@
  • Super-resolution filtering: Minimum Variance Distortionless Response (MVDR) [31, 33]
  • Spatial spectrum estimator: Multiple Signal Classification (MUSIC) [9, 11]
  • Compressed sensing: Sparse representations [29, 26]
  • Machine-learning based methods [28, 12, 16]
@@
The core differences between these listed techniques lie in how they approach the fundamental trade-off between image quality characteristics such as spatial resolution, sidelobe level, dynamic range, and robustness to noise as well as to correlated sources. Furthermore, computational cost has to be taken into account for practical, real-world systems. While advanced methods (like MVDR and MUSIC) offer superior resolution compared to conventional DAS, they often incur high computational complexity, and rely on assumptions that can degrade real-time performance and dynamic range in complex, real-world ultrasonic environments. More specifically, most of the so-called super-resolution techniques use spatial statistics of the incoming wavefield, requiring multiple measurements (snapshots) for the estimation thereof. In acoustic imaging, the luxury of multiple snapshots is not available, as this would incur a measurement time that is too large given the slow speed of sound[33].
@@
Given the constraints of real-time, high dynamic range acoustic imaging, there is a need for a spatial signal processing algorithm that offers better image quality than for example DAS, without incurring the computational complexity and model dependence of MVDR or MUSIC. In this paper, we take inspiration from medical ultrasonic imaging, a field where contrast and resolution are paramount, from their usage of non-linear beamforming methods, which have been shown to drastically enhance image contrast and suppress sidelobe levels. In the biomedical field non-linear beamforming has been extensively explored to enhance contrast and resolution. We focused on the application of the Delay-Multiply-and-Sum (DMAS) [14] combined with the optional Coherence Factor (CF) [5] weighting for in-air acoustic imaging. This methodology has demonstrated superior contrast ratios and increased range resolution in various biomedical applications, ranging from standard piezoelectric transducers to recent PMUT-based systems [24], but remains relatively unexplored in the context of in-air acoustic arrays.
The core principle of DMAS is to introduce a non-linear multiplication of delayed (or so-called pre-steered or pre-aligned) microphone signals. The DMAS algorithm as implemented for biomedical ultrasound by G. Matrone et al. [17] has seen various adaptations to reduce its computational load, such as the algebraic factorization proposed by A. Ramalli et al. [22], baseband implementations [25], and double-stage approximations [19]. While these optimization strategies have proven effective for standard second-order non-linearities, and specific higher-order formulations have been investigated for photoacoustic imaging [20], the specific requirements of in-air sonar necessitate an efficient, generalized approach capable of handling arbitrary orders.
@@
Figure 2: Directional component of the point-spread function in both azimuth and elevation, plotted using a Lambert Equal Area Projection. The resulting PSF shows a higher dynamic range for increasing orders of DMAS processing. Furthermore, the addition of the CF post-processing step further increases the dynamic range of the PSF. Gridlines are spaced by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$.
@@ II Methods
This section details the mathematical and computational methodology used to generate an acoustic image and perform source localization from raw microphone signals. The contribution of this paper focuses on the beamforming step of the imaging process, which will be reflected in this section as well. However, the overall process, as implemented in our general purpose signal processing pipeline[7], follows a sequential order with additional optional pre- and post-processing steps and filters, also visualized in Fig. 1.
These steps are:
@@
The system consists of an array of $N$ elements, in our case microphones. The received signal at the $i$-th microphone is denoted as $m_{i}(t)$. To form an image, a set of $M$ directions of interest is defined, which effectively constitutes the spatial grid of pixels or voxels where the acoustic intensity will be estimated. For each direction $\psi$, a time delay $\tau_{i,\psi}$ is calculated for each $i$-th microphone, representing the propagation time of an acoustic wave from the direction to the $i$-th microphone. In the case of 3D beamforming, these directions are defined by their azimuth and elevation angles $(\theta,\varphi)$. Fig. 1 visualizes the coordinate system and the signal processing pipeline.
The delays for these directions are pre-computed and stored in a delay matrix look-up table. The pre-steering operation is to time-shift the received signals according to these delays, effectively compensating them. This temporal alignment ensures that signal components originating from the target direction are in phase across all channels, a prerequisite for coherent summation. For a given time $t$ and direction $\psi$, we define the delayed signal $x_{i}(t,\psi)$ as:
@@ II-A Delay-and-Sum (DAS) Beamforming
DAS, as a non-adaptive beamformer, treats desired signals and noise/clutter identically, leading to high sidelobes and low image resolution and contrast, which is especially problematic in the noisy, attenuating environment of in-air acoustic imaging.
@@ II-B Delay-Multiply-and-Sum (DMAS) Beamforming
DMAS is a non-linear technique that involves multiplying combinations of delayed signals before summation [14, 17]. It is generally described by:
@@
The multiplication will cause lower sidelobes and a higher dynamic range by acting as a correlation function, which in effect suppresses low-coherence parts in the image signals. As these low-coherence parts are typically sidelobes in the image, and not real reflected signals, this operation increases image quality. High-order DMAS methods have been proposed to further suppress more coherent artifacts [20]. Let the $n$-th signed root of the delayed signal $x_{i}$ be:
@@
Using this $n$-th root signal $s_{i}^{(n)}$, the higher-order version of DMAS (order $n$) is given by:
@@ II-C Efficient Higher-Order DMAS Beamforming
To achieve an efficient implementation, we use the Newton-Girard identities [15], which can provide an explicit formula for $E_{n}$ in terms of the power sums of its arguments. The first step is to compute the necessary power sums, $P_{k}^{(n)}$, for $k=1,\ldots,n$:
@@
Note that order 2 is equivalent to DMAS as in Equation (3). Furthermore, the computational efficiency of our Newton-Girard implementation for the second order is mathematically equivalent to the factorization-based F-DMAS approach established by A. Ramalli et al [22]. However, unlike prior factorization or multi-stage methods [19] which are often derived specifically for quadratic or cubic non-linearities, the Newton-Girard identities provide a generalized, exact, and non-recursive formulation for any arbitrary order $n$. For any given order $n$, our method involves first computing the power sums $P_{1}^{(n)},\ldots,P_{n}^{(n)}$ in $\mathrm{O}(N)$ time, and then substituting these values into the corresponding explicit formula above to obtain $S_{\textsl{DMAS}}^{(n)}=E_{n}$. This non-recursive, explicit method is highly efficient for a fixed set of orders.
@@ II-D Coherence Factor (CF) Weighting
The Coherence Factor (CF) is an optional spatial weighting applied to the beamformer output to suppress incoherent noise and sidelobes [5]. It measures the spatial coherence of the delayed signals across the array aperture, and is defined as the ratio of the coherent power to the incoherent power. The specific formulation we implemented can be written as follows:
@@
II-E Computational GPU Acceleration and Code Availability
Even with the proposed efficient implementation, the entire signal processing is computationally demanding, especially for real-time processing on embedded or edge-computing devices. To achieve high performance, we have implemented the necessary beamforming operations on GPU hardware. We implemented a kernel function that achieves parallelism by launching a separate thread for each output pixel (i.e., for each $(t,\psi)$). The efficient $\mathrm{O}(N)$ DMAS formulation is critical, as it allows each thread to compute its output by iterating only once over the $N$ microphones. Our accelerated implementation is written in NVIDIA’s CUDA framework, and the source code can be found on our GitHub page: https://github.com/Cosys-Lab/Acoustic-Imaging-Toolbox. . This repository is structured as a Matlab Toolbox containing the proposed DMAS-CF methods implemented in native Matlab code, CPU-accelerated C, and GPU-accelerated CUDA.
@@ III Experimental Results And Validation
To evaluate the performance and merit of the DMAS algorithms for in-air sensing, we conducted several experiments. We want to note that in our results, we restrict our comparison solely to DAS against our proposed method because we want to focus on real-world conditions for a broad field of realistic application scenarios. This means we want to operate using only single snapshot, precluding the use of adaptive methods like MVDR which require multiple snapshots to accurately estimate data statistics. Furthermore, we require a method that functions effectively with any arbitrary microphone array configuration. Indeed, single-snapshot adaptations for super-resolution techniques exist, but require the application of spatial smoothing, which puts constraints on the array layout[33]. Finally, the imaging systems we want to apply these techniques to are inherently broadband, with typical frequency ranges from $25\text{\,}\mathrm{kHz}$ to $50\text{\,}\mathrm{kHz}$.
For these wideband applications, techniques like DAS and DMAS inherently function well as they are time-domain processing algorithms, unlike some frequency-domain or narrowband adaptive techniques such as MVDR or MUSIC.
@@ III-A Simulation Analysis
In order to quantify the performance of DMAS-class of beamformers in in-air sonar applications, we did several simulations of key metrics of this imaging system. As a first validation, we calculated the Point-Spread Function (PSF) of the whole imaging system. A PSF is a commonly used characterization of an imaging system, and describes the response of the system to a Dirac point source. In an active imaging system the point-spread function consists of three dimensions: two direction dimensions and one range dimension.
In Fig. 2 we have plotted the directional part of the PSF in the noise-free case. From this figure it becomes apparent that the dynamic range of the point-spread function becomes much greater for increasing DMAS orders, reaching almost $80\text{\,}\mathrm{dB}$ in the DMAS5 class. Compared to the approximated dynamic range of $40\text{\,}\mathrm{dB}$ of DAS, the improvement is substantial. Furthermore, the CF post-processing step further increases the dynamic range of the PSF. Little effect can be observed on the width of the main-lobe of the PSF. While the results in Fig. 2 show significant contrast enhancement, it should be noted that non-linear beamformers achieve this largely through dynamic range alteration rather than an increase in fundamental information content [23]. These methods non-linearly map the signal, which suppresses sidelobes but may not necessarily improve the underlying detectability of targets in all noise conditions. However, for in-air applications where visual localization and automated thresholding are key, this enhanced contrast remains a significant practical advantage.
A second view of the point-spread function can be made in the range/direction plane, more specifically in the horizontal plane for an elevation of $0\text{\,}\mathrm{\SIUnitSymbolDegree}$. Effectively, the PSF is sliced in the horizontal plane, and shown in a range/azimuth plot. The resulting ensemble of these PSFs can be found in Fig. 3. The dynamic range of the PSF improves for increasing orders of DMAS processing, with an additional gain in the dynamic range when applying the CF post-processing step. As expected, the range resolution is not impacted by the DMAS algorithms, because of the separability of the range and direction dimensions in far-field broadband imaging systems, highlighting the function of DMAS as a pure beamforming operation.
@@
The literature of medical imaging does highlight increased range resolution when applying DMAS, which can be explained by the fact that in medical imaging, the imaging takes place in the near-field regime (with the sensor array not being considered ”small” in respect to the imaging scene). Therefore, range-focusing through beamforming is possible, which causes an effect on the range resolution of the DMAS operations. As with in-air sonar sensing we typically operate in a pure far field regime (with array apertures of $10\text{\,}\mathrm{cm}$ diameter, and imaging ranges spanning up from $5\text{\,}\mathrm{m}$ to $10\text{\,}\mathrm{m}$), the separability of the range and directional imaging steps holds, and therefore, the lack of increase in range resolution when applying DMAS is expected.
In a second experiment, we calculated the signal-to-noise ratio (SNR) of the acoustic image for a varying SNR of the input microphone signals, using the following procedure: Let the amplitude of the echo signal for a source placed at broadside of the array at $1\text{\,}\mathrm{m}$ range in the microphone channels $m_{i}(t)$ be equal to 1. We add a white gaussian noise signal $n_{i}(t)$ to the microphone signals in the following way:
@@
We calculated these image SNR quantities for an input range of $\mathrm{mic}$ ranging from -$40\text{\,}\mathrm{dB}$ to $10\text{\,}\mathrm{dB}$. The results of these simulations can be found in Fig. 4, with the panels showing the SNR results for the imaging with and without coherence factor post-processing on the left and right respectively. From these plots the superior noise suppression of the DMAS-class of algorithms can be observed, with increasing image SNR values for increasing DMAS orders. The addition of the CF post-processing step further increases the image SNR over all algorithms. In Table I we have highlighted the results at a reference microphone SNR of $0\text{\,}\mathrm{dB}$ to further quantify the improvements in SNR by DMAS-CF.
TABLE I: Impact of CF and Beamforming Order on Image SNR at $0\text{\,}\mathrm{dB}$ Input
Algorithm No CF With CF CF Gain
SNR (dB) $\Delta$ (dB) SNR (dB) $\Delta$ (dB) $\Delta$ (dB)
DAS 32 55 +23
DMAS 46 14 66 11 +20
DMAS3 58 26 73 18 +15
DMAS4 65 33 76 21 +11
DMAS5 71 39 78 23 +7
As a final quantification experiment, we measured the spatial resolution of the DMAS-class beamformers. For this, we implemented a simulation in which two reflectors were spaced at the same range ($1.5\text{\,}\mathrm{m}$) and were placed symmetrically with respect to the X-axis with a half-angle $a$, as can be seen in Fig. 5 on the bottom right corner. We then calculated the acoustic image $I$ in the horizontal plane, and selected the range slice corresponding to the reflector range. We then aggregated the angular responses for all inner-angle combinations of reflectors in the scene, which are the results shown in Fig. 5. From these plots it becomes clear that the main benefit of the DMAS methods is an increase in dynamic range, but not in spatial resolution, which remains mostly constant over the spacing of the targets. This confirms the hypothesis that the spatial resolution mainly originates from the layout of the microphone array, and cannot be drastically improved by applying the DMAS-class algorithms for beamforming.
@@
To support the claim that the spatial resolution of the imaging system only marginally depends on the order of DMAS, we calculated the beamwidth of the various beamformer implementations for varying microphone array sizes. We simulated circular arrays with uniformly spaced microphones inside of the array, and varied the radius of the arrays from $1\text{\,}\mathrm{cm}$ to $6\text{\,}\mathrm{cm}$, for which the microphone count varied from 19 to 513 microphones (as we kept the spacing between the microphones constant on an hexagonal lattice with an edge length of $5\text{\,}\mathrm{mm}$). For each of the array shapes we then calculated the point-spread function in the horizontal plane and calculated the -$3\text{\,}\mathrm{dB}$ beamwidth of the PSF. This does not cause a loss of generality, as the PSF of a circularly symmetric array is circularly symmetric as well. We performed this calculation for all beamformer implementations, and plotted the resulting beamwidths as a function of the array radius in Fig. 6. The right panel details the experiment and illustrated the more narrow PSF as the array size increases. Both graphs show that the major decrease in beamwidth is caused by the array shape, and less so for the various beamformer implementations, with the beamwidths converging for the larger (yet still very practical) array sizes. This convergence of spatial resolution for larger arrays coincides with the findings depicted in Fig. 5, where it could be observed that the spatial resolution is not greatly increased using the DMAS-class of beamformers.